J. Sanikidze and K. Ninidze

N. Muskhelishvili Institute of Computational Mathematics, Georgian Academy of Sciences Tbilisi, Georgia

ON APPLICATION OF SINGULAR INTEGRAL EQUATIONS

TO A NUMERICAL SOLUTION OF SOME BOUNDARY PROBLEMS

Numerical solution of Dirichlet modified problem by singular integral equations is represented. Using the obtained computational schemes some concrete problems of conformal mapping are solved. The results of numerical experiments go well with theory.